You measure out 5.2 mL of an aqueous ethylenediamine (H2NCH2CH2NH2) solution. This aqueous solution has a density of 0.950 g/mL and is 25.0% by mass ethylenediamine. How many moles of ethylenediamine are in the 5.2 mL?
Answer #1

Solution-


We can find the mass of the solution by using = density*volume

=0.95*5.2

= 4.94 g

Now 100 gm of solution has 25 g of ethylenediamine
So 4.94 g of solution 25/100*4.94

=0.0506 g of ethlenediamine
As we know that the molar mass of ethylenediamine, MW = 60.1 g/mol
Moles of ethylenediamine, n = 0.0506/60.1

= 8.41*10^-4 mol